OpenStudy (anonymous):

A ball moving at +3.0 m/s along a table rolls off a table and lands on the ground 2.0 m away. How high was the table? Option A: 1.1m Option B: 2.2m Option C: 3.0m Option D: 3.3m I got 1.1m, but I'm not sure if that's correct? Thanksa yousa! :)

OpenStudy (anonymous):

At 3m/s the ball walked 2m, from the formula x=vt we get that t=2/3. We know that the initial velocity on the z axis and the initial position can be ignored, so: 1/2gt^2=h and h=20/9 and that is greater than 2 and lower than 3 so it is B.

OpenStudy (anonymous):

Ohhh alright, yeah, I was doing something totally different, haha. Your explanation really helped me understand this concept, so thanks again!